# Explore Synergistic Interaction Across Frames for Interactive Video Object Segmentation

### v1

@@ Abstract
Interactive Video Object Segmentation (iVOS) is a challenging task that requires real-time human-computer interaction. To improve the user experience, it is important to consider the user’s input habits, segmentation quality, running time and memory consumption. However, existing methods compromise user experience with single input mode and slow running speed. Specifically, these methods only allow the user to interact with one single frame, which limits the expression of the user’s intent. To overcome these limitations and better align with people’s usage habits, we propose a framework that can accept multiple frames simultaneously and explore synergistic interaction across frames (SIAF). Concretely, we designed the Across-Frame Interaction Module that enables users to annotate different objects freely on multiple frames. The AFI module will migrate scribble information among multiple interactive frames and generate multi-frame masks. Additionally, we employ the id-queried mechanism to process multiple objects in batches. Furthermore, for a more efficient propagation and lightweight model, we design a truncated re-propagation strategy to replace the previous multi-round fusion module, which employs an across-round memory that stores important interaction information. Our SwinB-SIAF achieves new state-of-the-art performance on DAVIS 2017 (89.6%, $\mathcal{J\&F}@60$). Moreover, our R50-SIAF is more than $\bf{3\times}$ faster than the state-of-the-art competitor under challenging multi-object scenarios.
@@
Grounded situation recognition, knowledge integration, trade-off weighting strategy
Figure 1: SIAF enables users to interact with multiple frames at once, whereas the SOTA method can only accept one single frame. Moreover, SIAF processes multiple objects simultaneously, leading to faster model performance and reduced computing resources.
@@ I Introduction
Previous methods have made significant progress. The coupling model [1, 2, 3, 4] jointly trains the interaction and propagation modules, but these coupling models might be trained hard. Then, the decoupled models [5, 6] were proposed, which train the interaction and propagation modules separately, such as the recent SOTA method that introduces a difference-aware fusion module [6].
@@
ii) ID-Queried Batch Processing. Previous methods suffer from slow running speed as they generate masks for each object individually. To enhance processing speed, a promising approach is to devise a model capable of handling multiple objects simultaneously. Nevertheless, a challenge lies in effectively distinguishing between these objects in a unified manner. To address this issue, we propose to assign a unique identifying tag to each object and train multiple distinct channels to efficiently process the associated scribbles and mask information linked to each tag. This id-queried batch method significantly reduces the running time from $F\times N$ to approximately $F$ when dealing with $N$ objects.
On the whole, we propose a framework that exploits Synergistic Interaction Across Frames (SIAF; Fig.2), which contains two main components: Across-Frame Interaction Module (AFI) and Across-Round Propagation Module. The AFI module generates masks for the interacted frames, while the Across-Round Propagation Module handles the remaining non-interactive frames. Concretely, we design the interactive module (see Fig.3) that can accept distinct scribbles on multiple frames and migrate scribbles information across frames, thus we can generate masks for all annotated objects and interacted frames based on the input scribbles. Moreover, during propagation, we propose an innovative across-round memory to store the important interacted information and update it for each round to generate masks. In addition, the truncated propagation strategy proposed can fully use across-round information by re-propagating for all frames.
Extensive experiments are conducted to validate the effectiveness. Particularly, our SwinB-SIAF achieves 89.6% $\mathcal{J\&F}@60$ on DAVIS after only three rounds of interaction. Meanwhile, our R50-SIAF runs more than $\bf{3\times}$ faster compared to the state-of-the-art method under challenging multi-object scenarios. Our code and benchmark will be released.
@@
  • We are the first to propose a framework called SIAF that enables users to annotate distinct objects on multiple frames in a single round, broadening the application scenarios and aligning better with people’s usage habits. Additionally, we introduced an interactive graphical user interface (GUI) that supports multi-frame interaction.
  • Our proposed Across-Frame Interaction Module is a novel approach that facilitates the transfer of scribble-based information between interactive frames. Moreover, we develop an Across-Round Propagation Module that can simultaneously process multiple objects in batches, and design a re-propagation strategy to effectively exploit the interactive information from multiple rounds.
  • We conduct extensive experiments on DAVIS 2017 [7] and the experimental results strongly demonstrate that our proposed SIAF runs $\bf{3\times}$ faster faster than the state-of-the-art method under challenging multi-object scenarios.
@@ II Related Work
II-A Semi-Supervised Video Object Segmentation
Semi-supervised video object segmentation involves segmenting a specific object with a fully-annotated mask in the initial frame. Early methods [6, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17] predominantly focus on modeling appearance representations and temporal connections. MoNet [18] fine-tunes pre-trained networks based on first-frame ground truth during testing, while OnAVOS [19] extends the first-frame fine-tuning by introducing an online adaptation mechanism. Similarly, MaskTrack [20] and PReM [21] utilize optical flow to propagate segmentation masks across frames. Nonetheless, these fine-tuning methods exhibit inefficiencies.
Recent advancements in video object segmentation methods have showcased innovative approaches [22, 23, 24, 25, 26]. For instance, PML [27] utilizes the nearest neighbor classifier to learn pixel-wise embedding, while VideoMatch [28] employs a soft matching layer, mapping current frame pixels to those in the initial frame. Building upon these foundations, FEELVOS [9] and CFBI [29, 30] extend the pixel-level matching mechanism by additionally matching between the current frame and the previous frame. STM [1] adopts a memory bank constructed from past frames and integrates a query key-value attention mechanism. Notably, AOT [31] introduces an identification mechanism to enable the simultaneous processing of multiple objects within a frame.
@@ II-B Interactive Video Object Segmentation (iVOS)
Interactive video object segmentation (iVOS) aims to produce real-time segmentation of a target object during user interactions, the interactions can be used to either segment an object or a correct previously misclassified region. Early methods use hand-crafted features (e.g.,click-based interactions [32, 33]), but they require a large number of user interactions, making them impractical at a large scale [34]. With the release of the DAVIS interactive benchmark [35], a standardization for the evaluation of iVOS has been introduced.
Previous approaches in user-guided segmentation often employed deep feature fusion techniques, utilizing interconnected encoder networks [36, 37, 25], or storing scribble features in memory for subsequent reference during segmentation [38, 4, 26]. However, these methods inherently link the specific user inputs with the mask propagation process, complicating model training. Balancing the understanding of user interactions and accurate mask propagation simultaneously poses challenges. To streamline the training process and enhance segmentation accuracy, recent methods aimed to decouple the interaction and propagation networks [39, 40]. This involves generating a mask based on various interaction types and then propagating this mask. For instance, IPN [1] and ATNet [2] use two segmentation networks to handle interaction and propagation, respectively. Cheng et al. [6] explore the differences in mask domains pre and post-interaction rounds, further refining the decoupling model.
@@
Figure 2: The overview of SIAF (§III-A). SIAF can be formulated as interaction-propagation. During interaction round $r$, the Across-Frame Interaction Module enables users to annotate multiple frames and generate the corresponding masks. Next, the Across-Round Propagation Module is designed to generate the masks for the remaining non-interactive frames. To preserve crucial interaction information, we update the across-round memory after each round. Then we generate masks with the updated memory by the batch decoder. We further introduce a re-propagation strategy to address conflicts in multi-round propagation.
@@ III-A Overview
Our pipeline for iVOS task can be formulated as interaction-propagation (depicted in Fig. 2). To address limitations in previous methods, such as single input mode and slow running speed, we carefully design two modules: the Across-Frame Interaction Module (AFI) and the Across-Round Propagation Module. These modules enable effective interactions across frames and facilitate information propagation across rounds.
Across-Frame Interaction Module (AFI). The Across-Frame Interaction Module (AFI; detailed in Sec.III-B) stands out from current methodologies reliant on single-frame inputs. In contrast, our AFI module excels in batch segmentation of multiple frames and diverse objects, showcasing its capacity to handle concurrent processing and enhance efficiency.
@@
where $AFI(\cdot)$ is the AFI module, which is utilized with scribbles, previous masks, and RGB images of multiple interacted frames, denoted as $\{{S_{t_{i}}^{r}},M_{{t_{i}}}^{r-1},{I_{t_{i}}}\}_{i=1}^{T}$, to generate output masks, denoted as $\{M_{t_{i}}^{r}\}_{i=1}^{T}$. If it is the first round of interaction, the previous masks are empty. Conversely, if it is a round following an interaction, the previous masks contain problems that require correction.
Across-Round Propagation Module The Across-Round Propagation Module aims to generate masks for the remaining non-interactive frames. In contrast to the previous propagation module that relied solely on the current round information and processes objects individually, our module leverages cross-round information and handles multiple objects in parallel. Concretely, we propose the Across-Round Propagation Module $Propagate(\cdot)$ to produce masks for non-interactive frames,
@@
Figure 3: The implementation of Across-Frame Interaction Module (AFI; §III-B). To capture the multi-scale features within each frame and the temporal information across frames, we design an across-frame attention (AFA). To enhance the modeling of the current frame, we develop the Feature Extraction Network. The premask at the top right of the image denotes the previous mask.
Figure 4: The attention map of Across-Frame Attention (AFA; §III-B). Continuing Fig. 3, showing the S2M in MiVOS that accepts only one single frame is insufficient. Suppose we aim to predict the mask of $t_{j}$-th frame. (d) The mask produced by S2M only refers to the $t_{j}$-th frame loss of the handlebar. To address this challenge, we utilize across-frame features to generate an attention map (b), which represents the handlebar’s attention position in the $t_{j}$-th frame (e) on the $t_{i}$-th frame (c). Thus, the mask (e) produced by SIAF contains the information of the other interacted frame (e.g. the handlebar information at $t_{i}$-th frame).
@@ III-B Across-Frame Interaction Module (AFI)
In contrast to existing methods that are restricted to single-frame and single-object scenarios, we propose the Across-Frame Interaction Module (AFI; Fig. 3), which is a novel method that can handle multiple frames and objects simultaneously. The AFI module comprises two sub-components: the Across-Frame Attention sub-component (AFA) and the Feature Extraction sub-component. The AFA sub-component captures temporal dependencies among frames, while the Feature Extraction sub-component extracts salient features from each frame.
@@
After obtaining the multi-scale features ${F}_{t}^{r}$, we divide them into two parts: ${\{\widetilde{F}_{t}^{r},{\hat{F}}_{t}^{r}}\}$. The middle-level feature, $\widetilde{F}_{t}^{r}$, captures across-frame information in the AFA sub-component, while the high-level and low-level feature, ${\hat{F}_{t}^{r}}$, extracts current-frame information in the Feature Extraction sub-component. This enables us to incorporate both the inter-frame and intra-frame information in our model.
Across-Frame Attention. To capture temporal information from multi-frame scribbles, we design a novel across-frame attention that integrates a deformable attention mechanism [41] and a multi-head attention mechanism [42]. The deformable attention allows for information exchange across different scales within each frame, while the multi-head attention facilitates information exchange between frames. This approach enhances the temporal coherence of the scribbles and enables effective modeling of objects.
@@
Feature Extraction. To improve feature utilization and model the current frame more effectively, we propose a Feature Extraction sub-component.
@@
where $\mathcal{G}(\cdot)$ denotes the Feature Extraction sub-component, which includes ASPP and FFN. ${\hat{F}_{t,curr}^{r}}$ is the feature embedding of current $t$ frame.
@@
Mask-Head Prediction. We implement a simple network as our mask head, incorporating Conv2D, BatchNorm, and Relu. The final mask $M_{t}^{r},{t}\in{R}$ is predicted using both the across-frame feature embedding ${\widetilde{F}}_{across}^{r}$ and the current frame feature embedding ${\hat{F}}_{t,curr}^{r}$ with the mask head.
@@
III-C Across-Round Propagation Module
Previous methods have primarily focused on modeling the current round while ignoring the interaction between different rounds. In light of this, we design the Across-Round Propagation Module, which maintains an across-round memory and exploits it for propagation. Moreover, we designed a batch decoder that can simultaneously process multi-object masks and images with identification tags in an efficient manner. The Across-Round Propagation Module is designed to generate masks for the non-interaction frames. To begin, we extract information from the interacted frame of the current round to obtain the interactive memory and employ it to update the across-round memory. Then we produce masks with the batch decoder by employing the across-round memory. In addition, we propose a truncated propagation strategy to address conflicts and an innovative re-propagation strategy to improve segmentation quality.

### v3

# IDPro: Flexible Interactive Video Object Segmentation by ID-queried Concurrent Propagation

@@ Abstract
Interactive Video Object Segmentation (iVOS) is inherently demanding, requiring real-time interaction between humans and computers. Enhancing user experience involves considerations such as user input habits, segmentation quality, running time, and memory consumption. However, existing methods compromise user experience by employing a single input mode and exhibiting slow running speeds. Specifically, these approaches restrict user interaction to a single frame, limiting the expression of user intent. To overcome these limitations and better align with user habits, we introduce a framework that facilitates flexible input modes by ID-queried concurrent propagation (IDPro). In particular, we have devised the Across-Frame Interaction Module (AFI), allowing users to freely annotate various objects across multiple frames. The AFI module transfers scribble information across interactive frames, generating multi-frame masks. Additionally, we leverage an id-queried mechanism to process multiple objects. To achieve more efficient propagation and a lightweight model, we propose a truncated re-propagation strategy, replacing the previous multi-round fusion module, which employs an across-round memory that stores crucial interaction information. Our SwinB-IDPro attains a new state-of-the-art performance on DAVIS 2017 (89.6%, $\mathcal{J\&F}@60$). Furthermore, our R50-IDPro exhibits over $\bf{3\times}$ faster performance than the leading competitor in challenging multi-object scenarios.
@@
Interactive Video Object Segmentation, Video Segmentation, Metric Learning, Identification Mechanism
Figure 1: IDPro enables users to interact with multiple frames at once, whereas the SOTA method can only accept one single frame. Moreover, IDPro processes multiple objects simultaneously, leading to faster model performance and reduced computing resources.
@@ I Introduction
Previous methods have made significant progress. The coupling model [1, 2, 3, 4] jointly trains the interaction and propagation modules, but these coupling models might be trained hard. Then, the decoupled models [5, 6, 7] were proposed, which train the interaction and propagation modules separately, such as the recent SOTA method that introduces a difference-aware fusion module [6].
@@
ii) ID-Queried Concurrent Propagation. Previous methods suffer from slow running speed as they generate masks for each object individually. To enhance processing speed, a promising approach is to devise a model capable of handling multiple objects simultaneously. Nevertheless, a challenge lies in effectively distinguishing between these objects in a unified manner. To address this issue, we propose to assign a unique identifying tag to each object and train multiple distinct channels to efficiently process the associated scribbles and mask information linked to each tag. This id-queried Concurrent method significantly reduces the running time from $F\times N$ to approximately $F$ when dealing with $N$ objects.
On the whole, we propose a flexible iVOS framework by ID-queried Concurrent Propagation (IDPro; Fig.2), which contains two main components: Across-Frame Interaction Module (AFI) and Concurrent Propagation Module. The AFI module generates masks for the interacted frames, while the Concurrent Propagation Module handles the remaining non-interactive frames. Concretely, we design the interactive module (see Fig.3) that can accept distinct scribbles on multiple frames and migrate scribbles information across frames, thus we can generate masks for all annotated objects and interacted frames based on the input scribbles. Moreover, during propagation, we propose an innovative across-round memory to store the important interacted information and update it for each round to generate masks. In addition, the truncated propagation strategy proposed can fully use across-round information by re-propagating for all frames.
Extensive experiments are conducted to validate the effectiveness. Particularly, our SwinB-IDPro achieves 89.6% $\mathcal{J\&F}@60$ on DAVIS after only three rounds of interaction. Meanwhile, our R50-IDPro runs more than $\bf{3\times}$ faster compared to the state-of-the-art method under challenging multi-object scenarios. Our code and benchmark will be released.
@@
  • We are the first to propose a framework called IDPro that enables users to annotate distinct objects on multiple frames in a single round, broadening the application scenarios and aligning better with people’s usage habits. Additionally, we introduced an interactive graphical user interface (GUI) that supports multi-frame interaction.
  • Our proposed Across-Frame Interaction Module is a novel approach that facilitates the transfer of scribble-based information between interactive frames. Moreover, we develop an Concurrent Propagation Module that can simultaneously process multiple objects, and design a re-propagation strategy to effectively exploit the interactive information from multiple rounds.
  • We conduct extensive experiments on DAVIS 2017 [8] and the experimental results strongly demonstrate that our proposed IDPro runs $\bf{3\times}$ faster faster than the state-of-the-art method under challenging multi-object scenarios.
@@ II Related Work
II-A Video Object Segmentation
The objective of Video Object Segmentation [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25] is to derive masks for target objects across the entire video. Methodologies commonly fall into the categories of semi-supervised VOS and unsupervised VOS. Unsupervised Video Object Segmentation (UVOS) specifically aims to autonomously distinguish primary foreground object(s) from the background in a video sequence. Existing methodologies have undergone extensive exploration. Notably, IMC-Net [26] introduces an implicit motion-compensated network that adeptly integrates complementary appearance and motion cues. In addition, [27] puts forth a contrastive motion clustering algorithm, while FEM-Net [28] introduces a motion-attentive encoder and a Flow Edge Connect module to infer edges in the optical flow for ambiguous or missing regions.
Another prevalent approach is semi-supervised VOS [29], semi-supervised video object segmentation involves segmenting a specific object with a fully-annotated mask in the initial frame. Early methods [30, 6, 31, 32, 33, 34, 35, 36, 37, 38, 39] predominantly focus on modeling appearance representations and temporal connections. MoNet [40] fine-tunes pre-trained networks based on first-frame ground truth during testing, while OnAVOS [41] extends the first-frame fine-tuning by introducing an online adaptation mechanism. Similarly, MaskTrack [33] and PReM [42] utilize optical flow to propagate segmentation masks across frames. Nonetheless, these fine-tuning methods exhibit inefficiencies.
Recent advancements in video object segmentation methods have showcased innovative approaches [43, 44, 45, 2, 3, 9]. For instance, [7] proposes an approach that considers pixel-wise similarities between reference and target frames, alongside the structural information of target objects. Similarly, OGS [46] introduces an innovative architecture based on object-aware global-local correspondence. PML [47] utilizes the nearest neighbor classifier to learn pixel-wise embedding, while VideoMatch [48] employs a soft matching layer, mapping current frame pixels to those in the initial frame. Building upon these foundations, CFBI [49, 50] extend the pixel-level matching mechanism by additionally matching between the current frame and the previous frame. STM [1] adopts a memory bank constructed from past frames and integrates a query key-value attention mechanism. Notably, AOT [51] introduces an identification mechanism to enable the simultaneous processing of multiple objects within a frame.
@@ II-B Interactive Video Object Segmentation (iVOS)
In contrast to other cross-modal video segmentation tasks like text-guided [52, 53, 54, 55, 56, 57, 58, 59] or audio-guided video segmentation [60, 61, 62, 63, 64], interactive video object segmentation (iVOS) methods [65, 66, 67, 68, 69] afford greater flexibility in incorporating user guidance information. Interactive video object segmentation aims to produce real-time segmentation of a target object during user interactions, the interactions can be used to either segment an object or correct previously misclassified region.
Early methods use hand-crafted features (e.g.,click-based interactions [70, 71]), but they require a large number of user interactions, making them impractical at a large scale [72]. With the release of the DAVIS interactive benchmark [73, 74], a standardization for the evaluation of iVOS has been introduced. Previous approaches in user-guided segmentation often employed deep feature fusion techniques, utilizing interconnected encoder networks [75, 76, 77, 2], or storing scribble features in memory for subsequent reference during segmentation [78, 4, 3]. However, these methods inherently link the specific user inputs with the mask propagation process [76], complicating model training. Balancing the understanding of user interactions and accurate mask propagation simultaneously poses challenges.
To optimize training procedures and improve segmentation precision, recent approaches have drawn inspiration from object tracking methodologies [39, 79, 80, 11, 12, 16, 81, 82]. These iVOS methods endeavor to disentangle interaction and propagation networks, as delineated in prior works [83, 84], which entails the creation of a mask through diverse interaction modalities, subsequently propagated for refined segmentation outcomes. For instance, IPN [1] and ATNet [2] use two segmentation networks to handle interaction and propagation, respectively. Cheng et al. [6] explore the differences in mask domains pre and post-interaction rounds, further refining the decoupling model.
@@
Figure 2: The overview of IDPro (§III-A). IDPro can be formulated as interaction-propagation. During interaction round $r$, the Across-Frame Interaction Module enables users to annotate multiple frames and generate the corresponding masks. Next, the Concurrent Propagation Module is designed to generate the masks for the non-interactive frames. To preserve crucial interaction information, we update the across-round memory after each round. Then we generate masks with the updated memory by the ID-queried decoder. We further introduce a re-propagation strategy to address conflicts in multi-round propagation.
@@ III-A Overview
Our pipeline for iVOS task can be formulated as interaction-propagation (depicted in Fig. 2). To address limitations in previous methods, such as single input mode and slow running speed, we carefully design two modules: the Across-Frame Interaction Module (AFI) and the Concurrent Propagation Module. These modules enable effective interactions across frames and facilitate information propagation across rounds.
Across-Frame Interaction Module (AFI). The Across-Frame Interaction Module (AFI; detailed in Sec.III-B) stands out from current methodologies reliant on single-frame inputs. In contrast, our AFI module excels in the segmentation of multiple frames and diverse objects, showcasing its capacity to handle concurrent propagation and enhance efficiency.
@@
where $AFI(\cdot)$ is the AFI module, which is utilized with scribbles, previous masks, and RGB images of multiple interacted frames, denoted as $\{{S_{t_{i}}^{r}},M_{{t_{i}}}^{r-1},{I_{t_{i}}}\}_{i=1}^{T}$, to generate output masks, denoted as $\{M_{t_{i}}^{r}\}_{i=1}^{T}$. If it is the first round of interaction, the previous masks are empty. Conversely, if it is a round following an interaction, the previous masks contain problems that require correction. This mask, derived from the preceding interaction and requiring refinement, is referred to as the imperfect mask.
Concurrent Propagation Module The Concurrent Propagation Module aims to generate masks for multiple non-interactive frames. In contrast to the previous propagation module that relied solely on the current round information and processes objects individually, our module leverages cross-round information and handles multiple objects in parallel. Concretely, we propose the Concurrent Propagation Module $Propagate(\cdot)$ to produce masks,
@@
Figure 3: The implementation of Across-Frame Interaction Module (AFI; §III-B). To capture the multi-scale features within each frame and the temporal information across frames, we design an across-frame attention (AFA). To enhance the modeling of the current frame, we develop the FrameEnhancer Network. The premask at the top right of the image denotes the previous mask.
Figure 4: The attention map of Across-Frame Attention (AFA; §III-B). Continuing Fig. 3, showing the S2M in MiVOS that accepts only one single frame is insufficient. Suppose we aim to predict the mask of $t_{j}$-th frame. (d) The mask produced by S2M only refers to the $t_{j}$-th frame loss of the handlebar. To address this challenge, we utilize across-frame features to generate an attention map (b), which represents the handlebar’s attention position in the $t_{j}$-th frame (e) on the $t_{i}$-th frame (c). Thus, the mask (e) produced by IDPro contains the information of the other interacted frame (e.g. the handlebar information at $t_{i}$-th frame).
@@ III-B Across-Frame Interaction Module (AFI)
In contrast to existing methods that are restricted to single-frame and single-object scenarios, we propose the Across-Frame Interaction Module (AFI; Fig. 3), which is a novel method that can handle multiple frames and objects simultaneously. The AFI module comprises two sub-components: the Across-Frame Attention sub-component (AFA) and the FrameEnhancer sub-component. The AFA sub-component captures temporal dependencies among frames, while the FrameEnhancer sub-component extracts salient features from each frame.
@@
After obtaining the multi-scale features ${F}_{t}^{r}$, we divide them into two parts: ${\{\widetilde{F}_{t}^{r},{\hat{F}}_{t}^{r}}\}$. The middle-level feature, $\widetilde{F}_{t}^{r}$, captures across-frame information in the AFA sub-component, while the high-level and low-level feature, ${\hat{F}_{t}^{r}}$, extracts current-frame information in the FrameEnhancer sub-component. This enables us to incorporate both the inter-frame and intra-frame information in our model.
Migrate Across-Frame Scribbles Information. To capture temporal information from multi-frame scribbles, we design a novel across-frame attention (AFA) that integrates a deformable attention mechanism [85] and a multi-head attention mechanism [86]. The AFA module is tasked with the simultaneous processing of information from multiple frames, necessitating considerable computational resources. In order to reduce the computational load while retaining the texture information of the image, we have chosen to use mid-level image features. Furthermore, since mid-level feature maps contain less information compared to low-level feature maps, and the attention mechanism can effectively capture global information in the image, we designed to use the attention mechanism to aggregate features over a larger range, thereby capturing the global relationships in the image. The deformable attention allows for information exchange across different scales within each frame, while the multi-head attention facilitates information exchange between frames. This approach enhances the temporal coherence of the scribbles and enables effective modeling of objects.
@@
Multi-Frame Masks Generation. To improve feature utilization and model the current interactive frame more effectively, we propose a FrameEnhancer sub-component. Concretely, low-level features contain a wealth of edge and texture information, while high-level features are rich in semantic information. Thus, we combine the low-level and the high-level features to better model the information of the current frame. Furthermore, convolution operations are efficient in processing high-resolution feature maps and can extract low-level features such as edges and textures from images through local receptive fields, they are more conducive to modeling the current frame. Therefore, we have applied convolution operations in the design of the FrameEnhancer sub-component.
@@
where $\mathcal{G}(\cdot)$ denotes the FrameEnhancer sub-component, which includes ASPP and FFN. ${\hat{F}_{t,curr}^{r}}$ is the feature embedding of current $t$ frame.
@@
We implement a simple network as our mask head, incorporating Conv2D, BatchNorm, and Relu. The final mask $M_{t}^{r},{t}\in{R}$ is predicted using both the across-frame feature embedding ${\widetilde{F}}_{across}^{r}$ and the current frame feature embedding ${\hat{F}}_{t,curr}^{r}$ with the mask head.
@@
III-C Concurrent Propagation Module
Previous methods have primarily focused on modeling the current round while ignoring the interaction between different rounds. In light of this, we design the Concurrent Propagation Module, which maintains an across-round memory and exploits it for propagation. Moreover, we designed an ID-queried decoder that can simultaneously process multi-object masks and images with identification tags in an efficient manner. The Concurrent Propagation Module is designed to generate masks for the non-interaction frames. To begin, we extract information from the interacted frame of the current round to obtain the interactive memory and employ it to update the across-round memory. Then we produce masks with the ID-queried decoder by employing the across-round memory. In addition, we propose a truncated propagation strategy to address conflicts and an innovative re-propagation strategy to improve segmentation quality.